Suppose you borrow Rs.800000 in student loan @8% p.a. the interest rate in 2019:-
(i) How much will you owe after 10 years of accumulating debt?
(ii) If inflation is 6% per year over this period, how much will you owe after 10 years in real 2019 Rupees?
(iii) Who is the loser in this scenario, you or the Bank?
1
Expert's answer
2020-07-08T17:48:42-0400

(i)"80 000\times(1+0.08)^{10}=172 714"


(ii) build-up multiplier

"(1+0.08)^{10}\times(1+0.06)^{10}=3.866"


accrued amount

"80 000\times3.866=309 304.47"


(iii)the second case is disadvantageous.

if the bank will provide credit conditions on the first occasion, then lost the bank.

if the bank provides credit conditions in the second case, then the bank won.